Please help me out on this problem.
Determine whether each ordered pair is a solution of the system
(1,-5)
4x-4=9
2x+3y=-13


Answer by bucky(2189)

To determine if the ordered pair is a solution to an equation, just plug the x and y values
of the ordered pair into the equation and see if the left side of the equation still
equals the right side.
So, is the ordered pair (1,-5) a solution to the equation 4x-4 = 9? Note that there is
no "y" term in the equation, so we do not need to be concerned with the value -5 in the
ordered pair. We only need to substitute the x value of the ordered pair into the equation.
So in place of the x in the equation we let x = 1 from the ordered pair. This makes the
equation:
4*(1) - 4 = 9
Simplifying this by doing the multiplication we get:
4 - 4 = 9
which further simplifies to 0 = 9. This obviously is not true. Therefore, the ordered
pair (1, -5) is not a solution to the equation 4x - 4 = 9
In the off chance that you meant the equation to be 4x - 4y = 9, if you substitute
1 for x and -5 for y the equation becomes:
4*(1) - 4*(-5) = 9
After the multiplications are done the equation is:
4 + 20 = 9
This is not true so the ordered pair (1, -5) is not a solution to the equation 4x - 4y = 9 either.
Next, is the ordered pair (1, -5) a solution to the equation 2x + 3y= -13. The same
process is used. From the ordered pair we get the values x = 1 and y = -5. Substitute
those values for x and y in the equation. When they are substituted the equation
becomes:
2*(1) + 3*(-5) = -13
By doing the multiplications this equation becomes:
2 - 15 = -13
which further simplifies to:
-13 = -13
This is true so the ordered point (1, -5) is a solution to the equation 2x + 3y = -13.
Hope these two examples are enough for you to see how to determine if an ordered pair is
a solution to an equation.